# FIT-GNN: Faster Inference Time for GNNs that ‘FIT’ in Memory Using Coarsening

### v4

@@ Appendix D Dataset Description
(a) Summary of datasets used for Graph classification Dataset
@@
(b) Summary of datasets used for Graph regression Dataset
@@
(c) Summary of datasets used for Node classification Dataset
@@
(d) Summary of datasets used for Node regression Dataset
@@ Appendix E Experiment Details (Parameters and Device Configuration)
For node classification and node regression tasks, we use Adam Optimizer with a learning rate of $0.01$ and $L_{2}$ regularization with $0.0005$ weight. We use Adam Optimizer with a learning rate of $0.0001$ and $L_{2}$ regularization with $0.0005$ weight for graph level tasks. For both coarsened graph (Gc-train) and set of subgraph based (Gs-train) model, we set epochs to 300 and the number of layers of GCN to 2 for training. We set the hidden dimensions for each layer of GCN to $512$.
The device configurations are Intel(R) Xeon(R) Gold 5120 CPU @ 2.202.20GHz, 256256GB RAM, NVIDIA A100 4040GB GPU. We use Pytorch Geometric to to train our models.
@@ Appendix F Results
Figure 5 presents the ablation study conducted to compare the performance of FIT-GNN for different experimental setup using different methods for appending nodes in the subgraph. The experiments are conducted on the Cora dataset for different coarsening ratios.
Figure 6 shows the time comparison for creating the subgraphs for Cora Dataset for different coarsening ratios. It is observed that the None method takes the least amount of time, which is obvious since the subgraphs are not appended with additional nodes. For the Extra Node Method, the time increases since the subgraph size increases as new nodes are appended. For Cluster Node, we observe the maximum time taken since not only are we looking at the neighboring nodes that are lost due to partition, but also we are keeping a track of the cluster to which these neighboring nodes belong, corresponding to which, we add the cross-cluster edges.
Table 10 presents the results on the node classification task with all the coarsening ratios $\{0.1,0.3,0.5,0.7\}$. To validate the use of “variation_neighborhoods” as a coarsening algorithm for all the results presented, we show an ablation study for different coarsening algorithms in Table 11 and 12. Table 13 shows the maximum GPU memory consumption during inference for different datasets with different coarsening ratios and methods of appending nodes. Note that this memory consumption is to store the graph and the weight matrix in the memory only.
@@
Cora Chameleon
kron $0.758\pm 0.004$ $0.800\pm 0.002$ $0.571\pm 0.004$ $0.580\pm 0.013$
PROTEINS ZINC (subset)
algebraic_JC $\mathbf{0.826}$ $\mathbf{0.739}$ $0.823$ $0.810$
kron $0.522$ $0.652$ $0.663$ $1.201$
heavy_edge $0.565$ $0.609$ $2.114$ $1.227$
variation_edges $0.652$ $0.565$ $1.548$ $1.883$
variation_cliques $0.696$ $0.652$ $0.689$ $0.742$

### v5

@@ Appendix D Dataset Description
(b) Summary of datasets used for Graph regression
@@
(c) Summary of datasets used for Node classification
@@
(d) Summary of datasets used for Node regression
@@ Appendix E Experiment Details (Parameters and Device Configuration)
For node classification and node regression tasks, we use Adam Optimizer with a learning rate of $0.01$ and $L_{2}$ regularization with $0.0005$ weight. We use Adam Optimizer with a learning rate of $0.0001$ and $L_{2}$ regularization with $0.0005$ weight for graph-level tasks. We set epochs to 20 and the number of layers of GCN to 2 for training. We set the hidden dimensions for each layer of GCN to $512$.
The device configurations are Intel(R) Xeon(R) Gold 5120 CPU @ 2.202.20GHz, 256256GB RAM, NVIDIA A100 4040GB GPU. We use Pytorch Geometric to train our models.
@@ Appendix F Results
Table 12 presents the results on the node classification task with all the coarsening ratios $\{0.1,0.3,0.5,0.7\}$. We did not compare our experiments with methods such as Cluster-GCN(Chiang et al., 2019) because the method focuses on reducing training time by sampling subgraphs for several iterations. During the inference phase, however, it still performs inference on the entire graph. Consequently, its inference performance would be at most comparable to that of a standard GCN.
Figure 6 shows the time comparison for creating the subgraphs for Cora Dataset for different coarsening ratios. It is observed that the None method takes the least amount of time, which is intuitive since the subgraphs are not appended with additional nodes. For the Extra Node Method, the time increases since the subgraph size increases as new nodes are appended. For Cluster Node, we observe the maximum time taken since we are not only looking at the neighboring nodes that are lost due to partition, but also are keeping a track of the cluster to which these neighboring nodes belong, corresponding to which, we add the cross-cluster edges.
@@
Table 13 shows the maximum GPU memory consumption during inference for different datasets with different coarsening ratios and methods of appending nodes. Note that this memory consumption is to store the graph and the weight matrix in the memory only.
@@
To validate the use of “variation_neighborhoods” as a coarsening algorithm for all the results presented, we show an ablation study for different coarsening algorithms in Table 14 and 15.
Cora Chameleon
kron $0.758\pm 0.004$ $0.800\pm 0.002$ $0.571\pm 0.004$ $0.580\pm 0.013$
PROTEINS ZINC (subset)
algebraic_JC $\mathbf{0.826}$ $\mathbf{0.739}$ $0.823$ $0.810$
kron $0.522$ $0.652$ $0.663$ $1.201$
heavy_edge $0.565$ $0.609$ $2.114$ $1.227$
variation_edges $0.652$ $0.565$ $1.548$ $1.883$
variation_cliques $0.696$ $0.652$ $0.689$ $0.742$
Appendix G Detailed Ablation Study on Node Regression Performance
To understand the counterintuitive performance gains observed in node regression tasks when using FIT-GNN, we conducted a comprehensive ablation study. Specifically, we isolated the source of the performance gap to determine why restricting inference to localized subgraphs outperforms full-graph inference.
G.1 Impact of Inference Input vs. Training Regime
Our first objective was to determine whether the performance gain stemmed from the training methodology itself or the structure of the input data (subgraphs) during inference. We evaluated two distinct setups on the Crocodile dataset using a GCN convolutional layer:
  • Setup A: Train on subgraphs $\rightarrow$ Infer on the full graph.
  • Setup B: Train on the full graph $\rightarrow$ Infer on the full graph.
Table 16: Comparison of training and inference setups to isolate the source of performance gains.
Train Setup Inference Setup Mean Absolute Error (MAE)
Full Graph Full Graph 0.852
Subgraphs Full Graph 0.865
Subgraphs (FIT-GNN) Subgraphs 0.364
As shown in Table 16, the performance of Setup A and Setup B is nearly identical, indicating that training on subgraphs alone does not account for the drastic MAE reduction. The substantial performance leap in FIT-GNN only occurs when subgraphs are used as the inference input, confirming that the local structural input drives the improvement.
G.2 Subgraph Optimization Landscape
We further hypothesized that partitioning the graph into subgraphs presents a simpler optimization landscape for the model. To quantify this, we compared the variation of labels locally (within individual subgraphs) versus globally across the entire graph. We computed the Label Standard Deviation for node regression datasets and Label Entropy for node classification datasets.
Table 17: Comparison of global label variation versus average local subgraph variation.
Dataset Metric Used Global Variation Subgraph Variation (Avg)
Cora Entropy 1.8311 0.1245
Citeseer Entropy 1.7533 0.1572
Chameleon Standard Deviation 2.1329 0.0689
Squirrel Standard Deviation 1.7639 0.1284
The results in Table 17 demonstrate that the variation of labels within individual subgraphs is drastically lower than the global variation. By coarsening the graph, we create local contexts that are statistically more homogeneous, allowing the GNN to specialize more effectively and navigate a smoother optimization landscape during inference.
G.3 Structural Information Loss and Implicit Adversarial Pruning
Finally, we investigated the role of structural information loss. Using a coarsening ratio of $r=0.5$, we empirically computed the fraction of the 2nd-hop neighborhood lost for each node.
(a) Citeseer (Classification)
(b) Cora (Classification)
(c) Squirrel (Regression)
(d) Chameleon (Regression)
Figure 7: Histograms detailing the fraction of the 2nd-hop neighborhood lost for each node at a coarsening ratio of $r=0.5$. A distinct difference in distribution is visible between classification (a, b) and regression (c, d) datasets.
Observations and Interpretation: The histograms in Figure 7 reveal a stark contrast between task types. In the node classification datasets (Cora, Citeseer), a significant number of nodes retain their 2nd-hop neighborhood entirely (lost fraction close to 0). Conversely, in the node regression datasets (Squirrel, Chameleon), the vast majority of nodes lose a highly significant portion of their 2nd-hop neighborhood (lost fraction close to 1).
Given the superior performance of FIT-GNN on these regression tasks, we interpret this structural loss as a form of implicit adversarial pruning. In specific heterophilic graphs, long-range information (such as 2nd-hop neighbors) acts as noise or an adversarial signal. The coarsening algorithms utilized by FIT-GNN implicitly filter out this distant noise. This pruning allows the model to fully exploit the low-variance, highly homogeneous local structures identified in Section G, thereby drastically reducing regression error.